OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through (5, -2) and slope=2/5 I'm on a really tight time crunch...if someone could give me a quick way to do problems like these i would be eternally grateful

OpenStudy (lalaly):

y=mx+c m is the slope and c is the y intercept so y=(2/5)x+c now u have to find c so u plug in the point u have (5,-2) in x and y -2=(2/5)(5)+c -2=2+c c=-4 so the equation is y=(2/5)x-4